# A Theory of Recommendations

### v2

@@ 5.1 Characterizing the Value of the Recommendation System
We now define the quantity
$\beta:=F(1/2-R).$
Due to Assumption 1, it has a simple interpretation: Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$. For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related. We then obtain the following result:
@@
Suppose Assumption 1 holds. Then, all types accept all recommendations and
$V(\beta)=\pi^{B}\Delta_{O}^{B}\text{.}$
To understand why all types accept all recommendations, the belief-updating decomposition in Section 4.1 is helpful. Its third component captures the potential shift of probabilities between controversial products, which induces a purely type-dependent payoff effect. Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$. Hence, a recommendation does not shift probabilities between the controversial products. As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result. [^20] An analogous argument applies for don’t-buy recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
5.2 Optimal Design of Recommendation Systems
We now use our characterization of the value of a recommendation system to address design issues, focusing on the optimal choice of the threshold. We start with the case of a symmetric population, then we treat asymmetric populations.
5.2.1 Symmetric Type Distribution
In the symmetric setting, we obtain a simple expression for $V(R)$, yielding transparent comparative statics results. [^21] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$. To this end, a reparametrization is helpful:
Definition 5.
Using these parameters, we immediately obtain:
Lemma 1.
Suppose Assumption 1 holds. Then,
$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
The value of the recommendation system depends on $\beta$ as well as on the prevalence of controversial products $Q$ and on the odds ratio $\sigma$ between good and bad products. Recall that, for each of the two controversial product versions, $\beta$ measures the fraction of the population willing to recommend it. Therefore, the parameter captures the influence of the distribution and the recommendation threshold on the value. [^22] Observe that by varying $R$ we can obtain any $\beta\in(0,1)$. We can interpret changes of $\beta$ as either changes of the threshold $R$ for a given distribution $F$ or as changes of the distribution $F$ for a given threshold $R$. The following result abstracts from this distinction, allowing for either interpretation. It shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
Lemma 2.
Suppose Assumption 1 holds. Then,
  • (i)
  • (ii)
The first statement in (i) holds because as $\beta$ increases a greater share of the population is willing to recommend a controversial product. Moreover, conditional on receiving a buy recommendation, the recommended product is thus more likely to be controversial. Therefore, the objective payoff component $\Delta_{O}^{B}$ decreases as $\beta$ increases, implying the second statement in (i). The effect of $\beta$ on $V(\beta)$ in (ii) reflects the effect on its two components. As $\sigma$ increases (good products become relatively more likely than bad ones), a buy recommendation becomes more likely. Hence, a high $\Delta_{O}^{B}$ is particularly valuable when $\sigma$ is high so that increasing $\sigma$ and lowering $\beta$ are complementary ($\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$). [^23] More technically, repeated application of the product rule shows that $\frac{\partial^{2}V}{\partial\beta\partial\sigma}=\frac{\partial\pi^{B}}{\partial\beta}\frac{\partial\Delta_{O}^{B}}{\partial\sigma}+\frac{\partial\pi^{B}}{\partial\sigma}\frac{\partial\Delta_{O}^{B}}{\partial\beta}+\frac{\partial^{2}\pi^{B}}{\partial\beta\partial\sigma}\Delta_{O}^{B}+\frac{\partial^{2}\Delta_{O}^{B}}{\partial\beta\partial\sigma}\pi^{B}$. The (incomplete) intuition just given corresponds to the first term being positive. The value-maximizing level of $\beta$ must thus be weakly decreasing in $\sigma$. The symmetry assumption turns out to render the problem linear in $\beta$, so the second result in (ii) follows. Using $\beta=F(1/2-R)$, Lemma 2 directly implies the following result. [^24] Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
@@ 5.2.1 Symmetric Type Distribution
By Lemma 2, an increase in $R$ reduces the probability of a buy recommendation, while it increases $\Delta_{O}^{B}$, the expected average payoff conditional on a buy recommendation. These effects exactly offset each other if and only if $\sigma=1$. Otherwise, one effect dominates the other. Based purely on uncontroversial products, a buy recommendation would be more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$). It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation. Put differently, the designer chooses between revealing whether the product is good or whether it is bad. He moves the threshold in the direction which yields full revelation with higher probability, which hinges only on the odds ratio of the uncontroversial product versions.
5.2.2 Asymmetric Type Distribution
For a symmetric population, the subjective component cancels out in the value of the recommendation system, allowing for a focus on the role of the objective content. Moreover, symmetry simplifies the analysis by ensuring that all types accept the recommendation. However, the assumption of a symmetric population may prove restrictive in some contexts, and it does not allow us to study the role of the subjective component for the optimal design of the recommendation system. To make headway in that direction, we introduce the following assumption.
@@ 5.2.2 Asymmetric Type Distribution
We thus capture asymmetry by means of a simple but flexible parameterization. By assuming that the controversial products are equally likely a priori, we ensure that any effect of a recommendation exclusively reflects the population asymmetry. Corollary 1 implies the following result:
Corollary 2.
Suppose Assumption 2 holds. Then, we have:
[…]
Thus, whether and how a recommendation shifts the probabilities of the controversial product versions depends entirely on the type distribution. When it is symmetric ($a=1$), there is no shift. When it is convex ($a>1$), so that more people prefer version $(1,0)$ to $(0,1)$, a buy recommendation shifts probability from $(0,1)$ to $(1,0)$, as it is more likely to come from a person who prefers $(1,0)$ to $(0,1)$. Similarly, when the type distribution is concave ($a<1$), the probability is shifted to $(0,1)$.
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept all recommendations depends on the threshold $R$. Reflecting this complication, there is no direct counterpart to Proposition 5. Nevertheless, the following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
@@
Figure 2: Illustration of Propositions 6 and 7
As illustrated in Figure 2, Proposition 6 provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). [^25] The figure might suggest that the parameter region where $R^{\ast}\rightarrow 1$ is in some sense larger than the area where $R^{\ast}\rightarrow 0$. This is an artefact of the otherwise quite convenient parameterization of the preference distribution but does not have any significant economic meaning. The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. We explore under which circumstances the conditions given in Proposition 6 are necessary as well as sufficient:
@@
Suppose Assumption 2 holds and fix $\sigma$. Then, as $Q\rightarrow 0$, $V(R)$ is
@@
    decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
@@
    increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
@@
    maximized for some $R^{\ast}\in(0,1)$ otherwise.
@@
The proof essentially proceeds in two steps. First, all types accept all recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.262626 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types accept all recommendations by Corollary 1. The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$. The proof shows that, even for larger values of $Q$, $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ often holds, explaining why the conclusion of Proposition 7 holds not only in the limit. This simplifies $V(R)$ by eliminating the case distinctions. Leveraging this, we can prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 6 are necessary, too.
The requirement that $Q\rightarrow 0$ is much stronger than necessary for this result. Even for higher values of $Q$, Figure 2 quite accurately separates the parameter range under which an interior solution exists from the area where it does not. [^27] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$. In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently close to 1. To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$. Hence, the value of the recommendation system is increasing in $R$ in the symmetric case, as the gain in the expected payoff conditional on receiving a buy recommendation outweighs the loss from the lower probability of a buy recommendation. This follows because the subjective effect in $V(R)=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ cancels out with symmetry. With an asymmetric population, the term $-\Delta_{S}^{B}\mathbb{E}[i]$ does not vanish. To illustrate this, consider a very asymmetric preference distribution such as $F(x)=(x+1/2)^{10}$, for which the probability of a type $i\leq 0$ is less than 0.1%. Then, $\mathbb{E}[i]>0$ and, by Corollary 2, $\Delta_{S}^{B}<0$ so that $-\Delta_{S}^{B}\mathbb{E}[i]>0$. Hence, the subjective effect is positive on average. Intuitively, though this population is distributed very asymmetrically around zero, it is essentially homogeneous in the sense that almost all types agree on the ranking of the controversial products: If such a product results in a buy recommendation, it almost surely has quality vector (1,0), as almost all types prefer this to $(0,1)$. As such, these products are not that controversial anymore, and even the subjective effect is “fairly objective.” Now, as $R$ increases, the probability of a buy recommendation still falls, and the “purely objective” payoff part $\Delta_{O}^{B}$ increases. However, for $R$ high enough, this “fairly objective” payoff part $-\Delta_{S}^{B}\mathbb{E}[i]$ also decreases, as an extreme threshold $R$ impedes the ability to distinguish between a good product and a widely preferred controversial product. While consumers are homogeneous regarding which controversial product version they prefer, they are still heterogeneous with respect to the intensity with which they do so. Hence, disentangling the good and the preferred controversial product is worthwhile. Therefore, an interior optimal threshold results.
5.3 Preference Heterogeneity and Controversial Products
As discussed after Proposition 7, preference heterogeneity only matters when controversial products exist. Moreover, the less homogeneous the population, the more likely it is that extreme thresholds are desirable so that receivers learn solely about uncontroversial products. Hence, one might conjecture that greater preference heterogeneity and a higher prevalence of controversial products reduce the value of recommendation systems. In the following, we address both issues in turn, demonstrating that these conjectures do not hold without qualification. For tractability, we focus on symmetric populations.
5.3.1 Preference Heterogeneity
In this section, we analyze how changes in the (exogenously given) preference distribution affect the value of the recommendation system. [^28] When we consider distinct sender and receiver distributions in Section 6.1, we will build further on this and take a design perspective instead of the comparative statics approach here. Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^29] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$. Further, for $R>1/2$ we must have $\beta\leq 1/2$ because $\beta=F(1/2-R)\leq F(0)=1/2$. Accordingly, for $R<1/2$ we must have $\beta\geq 1/2$. This corresponds to a mean-preserving spread of $F$ (see Figure 3), coinciding with an increase in $\beta$ for $R>1/2$ and a decrease for $R<1/2$. We obtain the following result.

### v3

@@ 4 Symmetric Populations
4.1 Value-Maximizing Recommendation Systems
Under Assumption 1, the value of the recommendation system can be expressed particularly concisely, as evidenced by the following result:
@@ 4.1 Value-Maximizing Recommendation Systems
If Assumption 1 holds, all types accept recommendations and
$V(R)=\pi^{B}\Delta_{O}^{B}\text{.}$
Intuitively, given the symmetry assumption, both controversial product versions lead to a buy recommendation with the same probability. Hence, a recommendation does not shift probabilities between the controversial products. As the other two (objective) effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, so that all types accept recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
We now use our characterization of the value of a recommendation system to address design issues, focusing on the optimal choice of the threshold.
@@
Intuitively, increasing $R$ makes the system more selective: fewer buy recommendations are issued, but those that are issued are more informative—i.e., they imply a higher expected payoff ($\Delta_{O}^{B}$). Conversely, lowering $R$ makes the system less selective, increasing the likelihood of buy recommendations at the cost of informativeness. When good products are more likely than bad ones ($\sigma>1$), it is optimal to be selective with buy recommendations, using a higher threshold to ensure that buy recommendations are more trustworthy. When bad products are more likely ($\sigma<1$), the opposite is true: being selective with don’t-buy recommendations is more valuable, and the platform should lower $R$. These forces exactly balance out when $\sigma=1$, so that the value of the system becomes independent of the threshold.
Put differently, the designer chooses whether the recommendation system should focus on identifying good or bad products. The threshold is set to reveal the more likely of these with greater certainty, aligning the selectiveness of the system with the underlying distribution of product quality.
4.2 Demand-Maximizing Recommendation Systems
We now turn to the analysis of demand-maximizing recommendation systems. Demand maximization requires a lenient recommendation threshold:
Proposition 6.
Suppose Assumption 1 holds. Then, demand is decreasing in $R$.
By Proposition 4 all types accept recommendations at all thresholds $R$ under Assumption 1. Thus, we can ignore the components of demand in equation (8) that come from receivers who do not accept don’t-buy recommendations. As a result, demand simplifies to $D(R)=\pi_{B}(R)$, and the result follows because the probability of buy recommendations is decreasing in $R$ [^16] See Lemma A.2 in the Appendix.
This result contrasts with Proposition 5 on value maximization. When $\sigma<1$, both value and demand are decreasing in $R$, so the objectives align. But when $\sigma>1$, they are diametrically opposed: while value increases with $R$, demand continues to fall. The underlying tradeoff is that increasing $R$ makes the system more selective—it raises the bar for a buy recommendation. From a value-maximization perspective, this can be desirable: fewer but more informative recommendations help receivers avoid bad matches, increasing their expected payoff. In this sense, selectivity enhances informativeness and thereby the value of the system.
By contrast, from a demand perspective, greater selectivity is never beneficial. The goal is simply to maximize the number of purchases, which is achieved by issuing buy recommendations as frequently as possible. A lenient system with a low threshold maximizes demand. Crucially, the symmetry assumption ensures that this leniency does not backfire: all types interpret recommendations in the same way and accept them. Thus, a low threshold increases demand without causing any mismatch between recommendation content and receiver preferences.
4.3 Preference Heterogeneity and Controversial Products
As discussed at the end of Section 3.1, preference heterogeneity only matters when controversial products exist. This naturally raises the question: does increasing heterogeneity—or increasing the share of controversial products—make recommendation systems less valuable? One might expect so, but the answer turns out to be more nuanced. To streamline the exposition, we focus only on value rather than demand in the two following subsections.
4.3.1 Preference Heterogeneity
To study the effect of changes in the preference heterogeneity, we consider mean-preserving spreads of the type distribution $F$, which preserve symmetry but increase the share of extreme types. It is useful to define the quantity $\beta:=F(1/2-R)$, which captures the fraction of the population willing to recommend a controversial product—either $(1,0)$ or $(0,1)$—at threshold $R$. For a fixed type distribution $F$, $\beta$ and $R$ are inversely related. Hence, we can interpret changes in $\beta$ as resulting either from changes in $R$ for a fixed distribution or from changes in the distribution itself at a fixed threshold. In what follows, we treat $\beta$ as a measure of polarization: for $R>1/2$, an increase in $\beta$ reflects a shift toward more extreme preferences, whereas for $R<1/2$, polarization corresponds to a decrease in $\beta$.
Proposition A.1 in Appendix A.2 shows formally how the value of the recommendation system depends on the threshold $R$ and the odds ratio $\sigma$. In the following, we sketch the main insights. Recall that $\beta$ represents the share of senders who are willing to recommend a controversial product. When $R>1/2$, a buy recommendation for a controversial product requires relatively strong preferences. [^17] An analogous argument applies to $R<1/2$. A higher $\beta$ then corresponds to an increase in polarization, making buy recommendations more frequent but less informative. Whether it is desirable to resolve this trade-off by increasing $\beta$ depends on $\sigma$: If good products are more likely than bad ones ($\sigma>1$), the system gains value from having buy recommendations that are highly informative, even if they are infrequent. In that case, polarization is harmful because it makes buy recommendations more common but more noisy. Conversely, when bad products are more likely ($\sigma<1$), the value of the system comes from being able to rule out bad products through lenient buy recommendations, and polarization helps by increasing their frequency.
4.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: all receivers would accept recommendations, as these would reveal whether the product is good or bad. Beyond this extreme case, a natural conjecture is that recommendations become less valuable as controversial products become more common. We find that this conjecture does not hold without qualification.
When good and bad products are similarly likely, a higher prevalence of controversial products indeed reduces the value of the recommendation system. But when one type of product is much more likely than the other, the value of the system may actually be maximized at an interior share of controversial products. A formal statement of this result is given in Corollary 4 in Appendix A.2.
The reason lies in the structure of the value expression, which is the product of the buying probability and the objective informativeness of a buy recommendation (see Proposition 4). As controversial products become more prevalent, good products become less likely, and fewer buy recommendations are triggered by objectively good products. At the same time, more buy recommendations stem from controversial products. Thus, both the frequency and the informativeness of buy recommendations change. Whether the overall value increases or decreases depends on which of these effects dominates and on the odds-ratio $\sigma$. When good products are relatively rare ($\sigma<1$), identifying and avoiding bad products becomes especially valuable, which can raise the value of the system despite a lower share of objectively good recommendations. Conversely, when bad products are rare ($\sigma>1$), the value can increase as the recommendation system becomes more effective at identifying good ones, even if some recommendations are noisy.
5 Asymmetric Populations
While providing a useful benchmark, the assumption of a symmetric population does not allow us to study the role of the subjective component for the optimal design of the recommendation system. To make headway in that direction, we study asymmetric populations in this section. For the sake of tractability, we make the following assumption.
@@ 5 Asymmetric Populations
We thus capture asymmetry by means of a simple but flexible parameterization. By assuming that the controversial products are equally likely a priori, we ensure that the subjective effect of a buy recommendation depends entirely on the asymmetry parameter $a$. When the type distribution is symmetric ($a=1$), there is no systematic bias in how controversial products are evaluated. As a result, buy recommendations do not shift beliefs about which controversial version the product might be—there is no subjective effect. In contrast, when the population is asymmetric (i.e., $a\neq 1$), one controversial version is systematically preferred over the other. In that case, buy recommendations tend to come from people who favor the majority-preferred version, and this introduces a systematic bias: The recommendation carries subjective information, shifting beliefs towards the more popular version.
5.1 Value-Maximizing Recommendation Systems
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept recommendations depends on the threshold $R$. The following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
@@ 5.1 Value-Maximizing Recommendation Systems
Figure 2: Illustration of Propositions 7 and 8
[…]
Proposition 7, illustrated in Figure 2, provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). [^18] The figure might suggest that the parameter region where $R^{\ast}\rightarrow 1$ is in some sense larger than the area where $R^{\ast}\rightarrow 0$. This is an artefact of the otherwise quite convenient parameterization of the preference distribution but does not have any significant economic meaning. The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. We explore under which circumstances the conditions given in Proposition 7 are not only sufficient but also necessary.
@@
Suppose Assumption 2 holds. If $Q$ is small enough, then
@@
    $V(R)$ is decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
@@
    $V(R)$ is increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
@@
    $V(R)$ is maximized for some $R^{\ast}\in(0,1)$ otherwise.
@@
The proof proceeds in two steps. First, all types accept recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation is dominated by the objective content. This simplifies $V(R)$ by eliminating the case distinctions. We can use this information to prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 7 are necessary, too.
The result clarifies when interior thresholds emerge despite the recommendation system being inherently coarse. In the symmetric case, where subjective preferences cancel out on average, value maximization requires choosing an extreme threshold—either very lenient or very selective—depending on which uncontroversial product is more likely. The goal is to make recommendations essentially objective. However, in an asymmetric population, recommendations systematically shift beliefs about controversial products in favor of the majority-preferred version. This subjective component can be informative, especially when controversial products are common and the population is largely aligned in which one they prefer. In that case, even a buy recommendation for a controversial product is not very controversial—it likely signals the version that most people want. As a result, it can be optimal to back off from extreme thresholds and instead choose an intermediate threshold that allows drawing inferences about the controversial products, too. When preferences are less aligned or controversial products are rare, this logic breaks down, and more extreme thresholds become optimal again.
We next discuss how the interior threshold $R^{\ast}$—when it exists—responds to changes in the key parameters of the model. [^19] Due to the intractability of the problem, all of these statements are based on numerical simulations, see Appendix B. First, $R^{\ast}$ is increasing in $\sigma$: as good products become more likely relative to bad ones, the recommendation system should be more selective, and a higher threshold becomes optimal. Second, the further the preference distribution departs from symmetry (i.e., the further $a$ is from 1), the more $R^{\ast}$ moves to the interior. Intuitively, as $a$ moves further away from 1, the population becomes more homogeneous in its ranking of controversial products, so that including subjective content into the recommendation is informative and not too controversial. Third, the effect of the prevalence of controversial products $Q$ is more subtle. As $Q$ decreases, the setting effectively approaches that of a symmetric population, since asymmetry becomes irrelevant in the absence of controversial products. Accordingly, $R^{\ast}$ moves closer to the extreme threshold that would be optimal in the symmetric case: it approaches 1 if $\sigma>1$ and 0 if $\sigma<1$.
5.2 Demand-Maximizing Recommendation Systems
Like value maximization, demand maximization becomes significantly more complex when the population is asymmetrically distributed. Nonetheless, we establish a sufficient condition under which a low recommendation threshold is optimal, mirroring the result in the symmetric case.